OpenStudy (anonymous):

Please help! Factor 64v^2-49u^2

OpenStudy (anonymous):

is this the answer (8v + 7u) * (8v - 7u)

OpenStudy (jdoe0001):

hmm how did you get that anyway?

OpenStudy (anonymous):

factor as a Difference of Squares

OpenStudy (jdoe0001):

ok.. yeap, you're correct

OpenStudy (anonymous):

so 64 is the square of 8, 49 is the square of 7, v2 is the square of v1, u2 is the square of u1 (8v + 7u) • (8v - 7u)

OpenStudy (jdoe0001):

\(\bf 64v^2-49u^2\quad \begin{cases} 64\to 8\cdot 8\to 8^2\\ \bf 49\to 7\cdot 7\to 7^2 \end{cases} \\ \quad \\ 64v^2-49u^2\implies (8^2v^2)-(7^2u^2)\implies (8v)^2-(7u)^2 \\ \quad \\ (8v-7u)(8v+7u)\)